OpenStudy (anonymous):

find the volume of the solid obtained by rotating about the y-axis the region bounded by the curve y=2x^2-x^3 and the x-axis. How do I do this?

OpenStudy (amistre64):

find the the x and y intercepts

OpenStudy (anonymous):

ok.

OpenStudy (amistre64):

y = x^2(2-3x) right?

OpenStudy (amistre64):

y=0 at 0 and 3/2 right?

OpenStudy (anonymous):

no y=x^2(2-x)

OpenStudy (amistre64):

thats right lol.... i transposed that 3 dint i :)

OpenStudy (amistre64):

so we have bounds of x=0 and x=2 then right?

OpenStudy (anonymous):

yes.

OpenStudy (amistre64):

we need to get these in terms of y to calculate the y axis spin; lets do an inverse

OpenStudy (anonymous):

ok sounds good.

OpenStudy (amistre64):

x = 2y^2 -y^3 ; solve for y :)

OpenStudy (amistre64):

wolfram might help with this :)

OpenStudy (anonymous):

good idea. I will get back to you a little later on this I have a baseball game.

OpenStudy (amistre64):

are you sure its spun around the y axis? and not the x axis?

OpenStudy (amistre64):

let me read up on the shell method; its prolly easier with this setup..

OpenStudy (amistre64):

2pi [S] x[f(x)] dx ; from [0,2]

OpenStudy (amistre64):

x(2x^2-x^3) = 2x^3 -x^4 soo... we integrate that to get; (5x^4 -2x^5)/10. At 0 it is zero; so we just worry about the 2

OpenStudy (amistre64):

i get 8/5 as the volume of the torus....if i did it right lol

OpenStudy (amistre64):

so i am going to try this on firefox..... and it seems to be working good so far :)